f(x) = kx * e ^ (- x / 2); k is a constant, x > 0 a) Find the value of k for f(x) to be a valid probability function


1
Expert's answer
2022-05-23T16:16:40-0400
"\displaystyle\int_{-\infin}^{\infin}f(x)dx=\displaystyle\int_{0}^{\infin}kxe^{-x/2}dx=1"

"\int xe^{-x/2}dx"

"u=x, du=dx"

"dv=e^{-x\/2}dx, v=-2e^{-x\/2}"

"\int xe^{-x/2}dx=-2xe^{-x/2}+2\int e^{-x/2}dx"

"=-2xe^{-x\/2}-4e^{-x\/2}+C"

"\displaystyle\int_{0}^{\infin}kxe^{-x/2}dx=k\lim\limits_{t\to \infin}\displaystyle\int_{0}^{t}xe^{-x/2}dx"

"=k\lim\limits_{t\to \infin}[-2xe^{-x/2}-4e^{-x/2}]\begin{matrix}
 t \\
 0
\end{matrix}"

"=k(-0-0-(-0-4))=4k=1"

"k=\dfrac{1}{4}"